A ball is thrown from a tower 42 m. high above the ground with a velocity of 270 m/sec directed at 280 from the horizontal. How long will the ball hit the ground?

1
Expert's answer
2022-04-19T15:27:40-0400

"v_y=v\sin 28°=127~\frac ms,"

"t_1=\frac {v_y} g=13~s,"

"h_1=\frac{v_y^2}{2g}=823~m,"

"t_2=\sqrt{\frac{2(h+h_1)}g}=13.3~s,"

"t=t_1+t_2=26.3~s."